Circular No. 7660

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 MD_{7} (LINEAR) M. Blythe, Lincoln Laboratory Experimental Test System, reports the

discovery by LINEAR of a comet on images obtained on July 11, when it appeared diffuse. Subsequent observations permitted identification with the object 2001 MD_{7}, so designated on MPS 31852 as a result of LINEAR observations made on June 21 and 24. L. Šarounová reports that CCD images of the comet obtained on July 12.9 UT at Ondřejov show a bright nucleus and faint coma.

2001 UT  α_{2000}  δ_{2000}  m_{2}
June 21.30593  19^{h}12^{m}54.^{s}22  −10^{o}23′10′′.6  17.6
July 11.21836  18 53 34.78  −12 20 04.6

Full astrometry is given on MPEC 2001-N27, as are the following orbital elements by B. G. Marsden (from 38 observations, June 21–July 12):

T = 2001 Nov. 30.4679 TT  ω = 244.^{o}4228
e = 0.670640  Ω = 129.3848 }2000.0
q = 1.261944 AU  i = 13.3635
a = 3.831499 AU n^{o} = 0.1314168 P = 7.500 years

MARS V. A. Krasnopolsky, Catholic University of America; and P. D. Feld-

man, Johns Hopkins University, write: “We obtained a spectrum (range 90–119 nm; spectral resolution 0.02 nm) of Mars on May 12 using the Far Ultraviolet Spectroscopic Explorer. Among other emissions, the spectrum reveals weak H_{2} lines of the Lyman-band system at 107.162, 111.861, and 116.676 nm. All of these lines originate from the v′ = 6, J′ = 0 level that is excited by the solar Lyman-β photons. The detected lines are the first observational evidence for molecular hydrogen on Mars. Detection of H_{2} on Mars was impossible for the numerous spacecraft that visited Mars; an attempt to detect H_{2} by averaging the Mariner 9 ultraviolet spectra resulted in an upper limit of 1.5 rayleighs for the H_{2} emission at 160.75 nm (Moos 1974, J. Geophys. Res. 79, 2887). H_{2} is a product of water-vapor chemistry on Mars, and the observed lines correspond to an H_{2} mixing ratio of ≈ 20 parts per million in the lower atmosphere.”

2001 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green